The lenght of a rectangle is 10 cm longer than its width. The peremeter of the rectangle is 68 cm. What are the length and width of the rectangle?
Answer by Alan3354(69443)

The lenght of a rectangle is 10 cm longer than its width. The peremeter of the rectangle is 68 cm. What are the length and width of the rectangle?
-------
Perimeter = 2W + 2L = 68
W + L = 34
L = W + 10
Sub for L in the 1st eqn
W + (W+10) = 34
2W + 10 = 34
W = 12
L = 22